# Steel-LLM: From Scratch to Open Source – A Personal Journey in Building a Chinese-Centric LLM

### v1

@@ 1 Introduction
Several initiatives, such as LLM360’s Amber (Liu et al., 2023; Tan et al., 2024), M-A-P’s Neo (Zhang et al., 2024b), and AI2’s OLMO series (Groeneveld et al., 2024a; OLMo et al., 2024; Muennighoff et al., 2024), have addressed these limitations by releasing complete training pipelines, datasets, and intermediate checkpoints. While these contributions are invaluable, their development typically demands extensive computational resources, making them inaccessible to smaller research teams or individual practitioners. This creates a significant gap for fully transparent and resource-efficient LLMs tailored for smaller-scale research efforts, particularly in non-English languages such as Chinese.
@@
Steel-LLM achieves competitive performance on benchmarks such as CEVAL and CMMLU, outperforming early models from larger institutions. The model architecture incorporates innovative features such as Soft Mixture of Experts (Soft MOE) and an enhanced Feed-Forward Network, optimizing performance within resource constraints. Our training framework, modified from TinyLlama (Zhang et al., 2024c), includes several optimizations for efficiency and usability, such as improved model loading, training progress restoration, and the ability to append data during the training process.
@@
  • Complete Transparency: We offer complete transparency in our development process, including the release of our training pipeline, dataset, model architecture, and intermediate checkpoints. This facilitates reproducibility and enables further research by the broader community.
@@ 2 Related Works
Recent developments in open-source LLMs have varied widely in terms of transparency and accessibility. Many models, such as Qwen (Bai et al., 2023a; Yang et al., 2024), Llama (Touvron et al., 2023b), Deepseek (DeepSeek-AI et al., 2024a), Gemma  (Team et al., 2024), InternLM  (Cai et al., 2024), Mixtral  (Jiang et al., 2023), Yi  (AI et al., 2024), GLM  (GLM et al., 2024), have been released with only the final model checkpoints and weights, while withholding crucial details such as training data, codes, and intermediate checkpoints. This limited transparency hinders reproducibility and makes it difficult for the broader research community to fully understand or build upon these models.
@@ 3.1 Soft MOE
Steel-LLM is trained using only on 8 GPUs with limited GPU memory. For a model featuring sparse MoE structure, all parameters must be loaded into the GPU memory. Owing to the sparse nature of expert selection, in contrast to a dense model, the FFN parameters of the sparse MoE model cannot be fully trained. Steel-LLM is a small-scale language model with 1 billion parameters, leading to a relatively low computational load. Meanwhile, our objective is to fully train each parameter and leverage the advantages of the MoE structure to enhance model performance. Consequently, we ultimately opt for the soft MoE(Puigcerver et al. (2024)) structure. Soft MoE is fully differentiable. Therefore, there is no need to consider problems such as expert unbalance that exist in sparse MoE.
@@ 3.2 Enhanced Feed-Forward Network
Since the Steel-LLM employs the soft MoE approach, within this framework, the Feed-Forward Network (FFN) effectively represents an expert. In a vanilla Transformer, The FFN comprises two layers of Multi-Layer Perceptrons (MLPs). In the architecture of large language models, a prevalent strategy is to enhance the first layer of the MLP using the SwiGLU(Shazeer (2020)) activation function(Bai et al. (2023a);Touvron et al. (2023a);DeepSeek-AI et al. (2024b)). The SwiGLU activation function enhances the model’s non-linear representational capabilities, thereby improving its performance. Additionally, we extended the application of the SwiGLU activation function to the second layer of the MLP within the FFN.
@@
The hyperparameters of Steel-LLM are presented in Table 1. We employed the Qwen1.5 tokenizer (Team (2024b)), which utilizes byte pair encoding (BPE).
@@ 4 Training Framework
To imporve training speed and conserve GPU memory, during the training process, we utilized techniques such as bfloat16 mixed-precision(Kalamkar et al. (2019)), Fully Sharded Data Parallel (FSDP, Zhao et al. (2023)), and FlashAttention(Dao et al. (2022)). Operator fusion represents an additional training optimization approach. By integrating multiple computational steps, it reduces intermediate activations and memory access, and can be implemented via CUDA or Triton. Specifically, during training, we adopted the CUDA-based version of Rotary Position Embedding (RoPE,Su et al. (2023)) and the Triton-based(Tillet et al. (2019)) cross entropy loss function. We then conducted an ablation study using a micro-batch size of 8 and a 1.8-billion-parameter model on NVIDIA A100 GPU to analyze the impact of these techniques on training efficiency and GPU memory usage, as presented in Table 2. Employing all training acceleration techniques, the training speed can be enhanced by approximately 50%.
@@ 5 Pretraining
To ensure consistency across the data, we initially standardized all datasets into a uniform format. Subsequently, we utilized Alibaba’s open-source tool, Data-Juicer, to meticulously filter and transcribe both the text and code data. We utilized a total of 21 text processing operators, as described in Appendix C, and 13 code processing operators, as listed in Appendix D. The final step involved employing the tokenizer from the Qwen1.5. This tokenizer was instrumental in converting the entire corpus into token-ids, which were then merged and segmented into manageable chunks. The distribution of the pretraining data across different domains is illustrated in Figure 2. Despite these efforts, our data preprocessing workflow still exhibits certain limitations, particularly the lack of balance in the proportion of texts from different domains within the pre-training corpus.
@@ Appendix A Pretraining Data Detailed Description
Zhihu QA Including 1 million entries, with 1.5GB in size.
BELLE (BELLEGroup, 2023) Train_2M_CN and train_3.5M_CN, generated by ChatGPT, contain 2 million and 3.5 million dialogue entries respectively, and were both used in this project. Note that these datasets are unverified and may contain errors.
@@
Firefly (Yang, 2023) Comprising 1.15 million entries covering 23 common Chinese NLP tasks and includes culturally relevant data such as couplets, poetry, classical Chinese translations, prose, and Jin Yong novels, resulting in a total of 1.15 million entries.
@@ Appendix B Supervised Finetuning Data Detailed Description
OpenHermes-2.5 (Teknium, 2023) Consisting of samples synthesized by large models and chat samples, filtered from open-source data like Airoboros, ChatBot Arena, and Evol Instruct, totaling 1 million entries.

### v2

@@ 1 Introduction
Several initiatives, such as LLM360’s Amber (Liu et al., 2023; Tan et al., 2024), M-A-P’s Neo (Zhang et al., 2024b), and AI2’s OLMO series (Groeneveld et al., 2024a; OLMo et al., 2024; Muennighoff et al., 2024), have addressed these limitations by releasing complete training pipelines, datasets, and intermediate checkpoints. While these contributions are invaluable, their development typically requires extensive computational resources, making them inaccessible to smaller research teams or individual practitioners. This creates a significant gap for fully transparent and resource-efficient LLMs tailored for smaller-scale research efforts, particularly in non-English languages such as Chinese.
@@
Steel-LLM achieves competitive performance on benchmarks such as CEVAL and CMMLU, outperforming early models from larger institutions. The model architecture incorporates innovative features such as Soft Mixture of Experts (Soft MoE) and an enhanced Feed-Forward Network, optimizing performance within resource constraints. Our training framework, modified from TinyLlama (Zhang et al., 2024c), includes several optimizations for efficiency and usability, such as improved model loading, restoration of training progress, and the ability to append data during the training process.
@@
  • Complete Transparency: We offer complete transparency in our development process, including the release of our training pipeline, dataset, model architecture, and intermediate checkpoints. This facilitates reproducibility and allows for further research by the broader community.
@@ 2 Related Works
Recent developments in open source LLM have varied widely in terms of transparency and accessibility. Many models, such as Qwen (Bai et al., 2023a; Yang et al., 2024), Llama (Touvron et al., 2023b), Deepseek (DeepSeek-AI et al., 2024a), Gemma  (Team et al., 2024), InternLM  (Cai et al., 2024), Mixtral  (Jiang et al., 2023), Yi  (AI et al., 2024), GLM  (GLM et al., 2024), have been released with only the final model checkpoints and weights, while withholding crucial details such as training data, codes, and intermediate checkpoints. This limited transparency hinders reproducibility and makes it difficult for the broader research community to fully understand or build upon these models.
@@ 3.1 Soft MoE
Steel-LLM is trained using only 8 GPUs with limited GPU memory. For a model featuring sparse MoE structure, all parameters must be loaded into the GPU memory. Owing to the sparse nature of expert selection, in contrast to a dense model, the FFN parameters of the sparse MoE model cannot be fully trained. Steel-LLM is a small-scale language model with 1 billion parameters, leading to a relatively low computational load. Meanwhile, our objective is to fully train each parameter and leverage the advantages of the MoE structure to enhance model performance. Consequently, we ultimately opt for the soft MoE(Puigcerver et al. (2024)) structure. Soft MoE is fully differentiable. Therefore, there is no need to consider problems such as expert imbalance that exist in sparse MoE.
@@ 3.2 Enhanced Feed-Forward Network
Since Steel-LLM employs the soft MoE approach, within this framework, the Feed-Forward Network (FFN) effectively represents an expert. In a vanilla Transformer, The FFN comprises two layers of Multi-Layer Perceptrons (MLPs). In the architecture of large language models, a prevalent strategy is to enhance the first layer of the MLP using the SwiGLU(Shazeer (2020)) activation function(Bai et al. (2023a);Touvron et al. (2023a);DeepSeek-AI et al. (2024b)). The SwiGLU activation function enhances the model’s non-linear representational capabilities, thereby improving its performance. Additionally, we extended the application of the SwiGLU activation function to the second layer of the MLP within the FFN.
@@
The hyperparameters of Steel-LLM are presented in Table 1. We employ the Qwen1.5 tokenizer (Team (2024b)), which utilizes byte pair encoding (BPE).
@@ 4 Training Framework
To improve training speed and conserve GPU memory, during the training process, we utilized techniques such as bfloat16 mixed-precision(Kalamkar et al. (2019)), Fully Sharded Data Parallel (FSDP, Zhao et al. (2023)), and FlashAttention(Dao et al. (2022)). Operator fusion represents an additional training optimization approach. By integrating multiple computational steps, it reduces intermediate activations and memory access, and can be implemented via CUDA or Triton. Specifically, during training, we adopted the CUDA-based version of Rotary Position Embedding (RoPE,Su et al. (2023)) and the Triton-based(Tillet et al. (2019)) cross entropy loss function. We then conducted an ablation study using a micro-batch size of 8 and a 1.8-billion-parameter model on NVIDIA A100 GPU to analyze the impact of these techniques on training efficiency and GPU memory usage, as presented in Table 2. Employing all training acceleration techniques, the training speed can be enhanced by approximately 50%.
@@ 5 Pretraining
To ensure consistency across the data, we initially standardized all datasets into a uniform format. Subsequently, we utilized Alibaba’s open-source tool, Data-Juicer, to meticulously filter and transcribe both the text and code data. We utilized a total of 21 text processing operators, as described in Appendix C, and 13 code processing operators, as listed in Appendix D. The final step involved employing the tokenizer from Qwen1.5. This tokenizer was instrumental in converting the entire corpus into token-ids, which were then merged and segmented into manageable chunks. The distribution of the pretraining data across different domains is illustrated in Figure 2. Despite these efforts, our data preprocessing workflow still exhibits certain limitations, particularly the lack of balance in the proportion of texts from different domains within the pre-training corpus.
@@ Appendix A Pretraining Data Detailed Description
Zhihu QA Including 1 million entries of questions and answers, with 1.5GB in size.
BELLE (BELLEGroup, 2023) Including train_2M_CN and train_3.5M_CN, which are generated by ChatGPT, containing 2 million and 3.5 million dialogue entries respectively, and both used in this project. Note that these datasets are unverified and may contain errors.
@@
Firefly (Yang, 2023) Comprising 1.15 million entries which cover 23 common Chinese NLP tasks and include culturally relevant data such as couplets, poetry, classical Chinese translations, prose, and Jin Yong’s novels, resulting in a total of 1.15 million entries.
@@ Appendix B Supervised Finetuning Data Detailed Description
OpenHermes-2.5 (Teknium, 2023) Consisting of samples synthesized by large models and chat data, filtered from open-source data like Airoboros, ChatBot Arena, and Evol Instruct, totaling 1 million entries.